The magnetic flux density at the centre 100 loop circular coil of 15cm diameter can be the same as to the magnetic flux density at the end of a 200 turns 500cm long solenoid with a current of 3A. estimate the magnitude of current that must flow through circular coil for this to happen (Take permeability of the free space as 4π × 10–7 H/M
1
Expert's answer
2020-05-29T09:47:06-0400

Magnetic flux density is measured by "\mu_o \frac NlI"

equating for both we get

"\frac{100}{15}\times I=\frac{200}{500}\times3"


"I=0.18 A"